OpenStudy (anonymous):

Prove that any prime of the form 3m+1 for some m in the set of natural numbers is also of the form 6n+1 for some n in the set of natural numbers

OpenStudy (anonymous):

if it is prime of the form \[3m+1\] then it is odd, so it is also of the form \[2k+1\]

OpenStudy (anonymous):

and then 2k+1 = 6m.... ? would that be the next part?

OpenStudy (anonymous):

does that help? because it tells you that \[3m+1=2k+1\] so \[3m=2k\]

OpenStudy (anonymous):

say basically that m is even

OpenStudy (anonymous):

if you want to be more formal say that since \[3m=2k\] you must have \[2| m\] since clearly 2 does not divide 3. therefore \[m=2n\] for some n and so \[3m+1=6n+1\]

OpenStudy (anonymous):

think that is as formal as you can be in this problem

OpenStudy (anonymous):

Oh ok, I see. Thank you, I always have trouble knowing how to start a proof.